# Aging-aware Energy Management for Residential Multi-Carrier Energy Systems

### v2

@@ 3 Policy design
where $\eta$ denotes a conversion factor or efficiency, $Q_{\textrm{TESS}}$ is the capacity in kWh. The thermal balance comes in as:
The electric power balance, on the other hand, is
@@ 3.1 Batteries
The remaining devices in the MCES are all battery-based ESS. Batteries have complex nonlinear dynamics, and several modeling techniques are presented in the literature [17]. In this work, models coming from empirical and physics-based approaches are used. The modeling is divided into two different sub-models: performance and aging. Under the UMF, this is represented in the transition function $\tilde{S}^{M}_{b,t}(\tilde{S}_{b,t},x_{b,t}|\theta_{b,t})$, which contains both the perf. model $p_{b,t}^{M}(.)$ and the aging model $d_{b,t}^{M}(.)$. The performance model predicts stored energy $SoC_{b,t}$ and terminal voltage $v_{t,b,t}$. The aging model is used to update the parameters of $p^{M}_{b,t}(.)$, as shown in Fig. 3. Even though the change in parameters $\theta_{b,t}$ becomes significant after considerable ageing has occurred, optimizing it in the short term can lead to considerable savings in the long and medium term [39, 40, 15]. This is because in PB ageing models the relationship between ageing states and controls actions is explicit and the policy can directly dynamically minimize it through the internal states such as $SoC_{b,t}$.
@@
The aging model $d_{b,t}^{M}(.)$ is a set of equations that describe the dynamics of the performance parameters $\theta_{b,t}$.
@@
$SoC_{\textrm{BESS},t^{\prime}_{1}}=SoC_{\textrm{BESS},t^{\prime}_{1}+\frac{H}{2}}$ (22)
where $t_{0}\leq t_{1}\leq H\leq T$. This way the OCP is better conditioned but the planner still has the freedom to decide $SoC_{\textrm{BESS},24\textrm{hs}}$. The proposed terminal condition has two key properties: is more flexible than fixing $SoC_{\text{BESS},24\text{hs}}=50\%$ and it bounds the value function $V_{H}$ of the OCP. Ideally, no terminal condition would be used to freely use all 3 storage systems. Unfortunately, to solve such an unbounded OCP, an optimization horizon $H$ much larger than 48 hours would be required [41, 42].
@@ 3.1.1 Performance models $p_{b,t}^{M}$
For the performance submodel, two alternatives have been implemented: a simple bucket model (BM) and a first order equivalent circuit model (ECM). A basic BM of the operation of a battery assumes that its output voltage $v_{t}$ is linear with the state of charge $SoC$, assuming no voltage drop. Hence the only equations of this model are:
$SoC_{b,t+1}=SoC_{b,t}-\frac{\Delta t}{Q_{b,t}.3600}.\eta_{c}.i_{b,t}\,,$ (23)
$i_{b,t}=\frac{P_{b,t}}{v_{t,b,t}.N_{s,b}.N_{p,b}}\,,$ (24)
$OCV_{b,t}=a_{OCV,b}+b_{OCV,b}.SoC_{b,t}\,,$ (25)
$v_{t,b,t}=OCV_{b,t}\,,$ (26)
$S_{b,t}=[SoC_{b},v_{t,b},i_{b}]^{T}_{t}$ (27)
where $i_{b,t}$ is the current passing through the cell, $OCV_{b,t}$ is the open circuit voltage, $\eta_{c}$ is the Coulombic efficiency [18] and $Q_{b,t}$ is the cell capacity in Ah. Each battery pack is assumed to be organized as a series connected module (SCM) where $N_{s/p,\ b}$ are the series cells per branch and parallel branches, respectively. In this model, the most relevant parameter in $\theta_{b}$ is the $Q_{b}$.
Figure 4: First order Equivalent Circuit Model.
A first-order ECM has improved accuracy due to the incorporation of diffusion and series resistance, Fig. 4. The performance sub-model $p_{b,t}^{M}(.)$ is then modified by adding the equation:
@@
where $i_{R_{1},b,t}$ is the current flowing through $R_{1}$ in Fig. 4. Eqs. 23, 24 and 25 are maintained. The ECM incorporates the series voltage drop that limits power output and the first-order diffusion dynamics. Here the relevant parameters are $\theta=[Q,R_{0}]^{T}$ which usually define the cell’s state of health $SoH$.
@@
For the aging models, the first alternative is an empirical sub-model presented by [28]. The empirical sub-model reduces all the degradation mechanisms into calendar and cyclic aging.
@@
where the fitting parameters $c_{1,...,5}$ are taken from [43] and $t_{0,b}$ is the elapsed lifetime of the battery $b$.
For the physics-based alternative, the reduced order model (PBROM) from [15] is used. It accounts for two degradation mechanisms: the solid electrolyte interface (SEI) and active material loss (AM). The author also presents a PBROM for Li-plating but its particular functional form prevents it from being implemented inside an NLP.
@@
$\eta_{k,b,t}=\frac{2.R.T}{F}.\text{sinh}^{-1}\left(\frac{i_{b,t}}{n_{\textrm{SEI}}.a_{s}.A.L_{n}.i_{0}}\right)$ (33)
$z_{b,t}=SoC_{b,t}.(z_{100\%,b}-z_{0\%})+z_{0\%,b}$ (34)
$\beta_{b}=e^{\frac{n_{\textrm{SEI}}.F}{R.T}.\left(\eta_{k,b}+OCV_{n,b,t}-OCV_{s,b}\right)}$ (35)
where $\eta_{k}$ is the SEI side reaction kinetic overpotential, $z$ is the Li stochiometry of the cell, $OCV_{n}$ is the open-circuit voltage of the anode made with an empirical fit, $OCV_{s}=0.4V$ is the side reaction open-circuit voltage, and $T$ is the cell temperature. It is assumed that the temperature $T$ is constant over time and is controlled by the local primary control system. For the BESS this assumption could be challenged given that the BESS is owned/operated by the same house owners. In the case of the EV, the ownership of the car battery pack sometimes correspond to the OEM/car manufacturer. Moreover not even residential energy flexibility oriented protocols such as S2 [44] allow the control of internal temperature of the storage assets. The rest of the parameters can be found in the A.
@@
The total aging is the contribution of both mechanisms SEI layer growth and AM loss. The capacity fade current is:
@@
which is later used again in 31.
@@
To model the power fade (i.e. the increase of $R_{0}$), the SEI layer thickness $\delta_{\textrm{SEI},b,t}$ growth is described by:
@@
Hence the dynamics of the series resistance $R_{0}$ are:
@@
The solvent S leaves the electrolyte to form the SEI layer thus, the volume fraction of S evolves with:
@@ 3.2 Electric Vehicle
where $P_{\text{tot},\text{EV},t}$ is the total power of the EV, $P_{\text{EV},t}$ is the charger power, and $P_{\text{drive},\text{EV}}$ is the power consumed driving assuming no public charging. The total power $P_{\text{tot},\text{EV},t}$ is then used in Eq. (24) and later for calculating the aging of the EV batteries. The average driving power is also sampled from a Gaussian distribution $P_{\text{drive},\text{EV}}\sim\mathcal{N}(\mu_{\text{drive}},\,\sigma_{\text{drive}}^{2})$. This is because the EV battery pack degradation during driving needs to be accounted for in the operation strategy (charging and driving)
@@
$\varepsilon_{SoC}=SoC_{\textrm{EV}}(t_{\textrm{dep}})-SoC_{\textrm{dep}}^{*}$ (44)
where $w_{SoC}$ is a weight chosen by the user.
@@
1:Define setpoint sequences $\mathcal{P}_{a,[t,t+(H+\Delta t)/2]}$
2:Define exogenous information sequences $\mathcal{W}_{[t+1,t+1+(H+\Delta t)/2]}$
3:Recalculate $P_{\textrm{TESS},t}$ and $P_{\textrm{grid},t}$ with Eqs. 18 , 19, $\mathcal{P}_{a,[t,t+(H+\Delta t)/2]}$, and $\mathcal{W}_{a,[t+1,t+1+(H+\Delta t)/2]}$
@@
The simulator is used to evaluate the policy $\pi$ and close the loop with the state measurements. It is designed to:
@@
The whole process is defined in Algorithm 1. First, the power setpoints must be adjusted for the grid and TESS because the forecast used in $X_{t}^{\pi}$ will never be the same as the actual exogenous inputs. Take a look at the balances, Eq. 18 and 19, which contain the loads and solar generation. It is clear that $\tilde{P}\neq P$ and a device must compensate for that difference. Thus, the simulator $S_{a,t}^{M}$ recalculates:
@@
Second, once these powers have been adjusted, the simulator uses them to obtain the true/actual/fidelity state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$. For the TESS it recalculates Eq. 17. For the $b$, it uses LiiBRA.jl [46] to swiftly simulate PBROM s of the performance of the battery [17, 47]. After that, the models from Jin [15] are used to calculate the true degradation outcome of the decisions $P_{a,t}^{*}$. Again, the reader must remember that the capacity fade (decrease in $Q_{b,t}$) is modeled in both the simulator $S^{M}_{a,t}$ and the approximate model of the planner $\tilde{S}_{a,t}^{M}$, whereas the power fade (increase in $R_{0,as,t}$) is only addressed in the simulator $S^{M}_{a,t}$. Finally, if an action $P_{b,t}^{*}$ causes a future state to go out of bounds ($S_{b,t+1}\leq\underline{S}_{b}$ or $S_{b,t+1}\geq\overline{S}_{b}$), the remaining actions are rejected and the $b$ remains in that state (either $\overline{S}_{b}$ or $\underline{S}_{b}$) until the next day. Finally, the carriers are re-balanced if necessary.
@@ 5 Case studies
5:     Simulate $S_{a,t+1}=S_{a,t}^{M}(S_{a,t},P^{*}_{a,t},W_{t+1})\ \forall t\in\mathcal{D}_{t}=[t;t+\frac{(H+\Delta t)}{2}]$, using Algorithm 1
@@
7:     Move time window $\mathcal{D}^{\pi}_{t}\leftarrow\mathcal{D}^{\pi}_{t}+\frac{(H+\Delta t)}{2}$;
@@
The building has a grid connection with a smart meter with 15min resolution. The connection is also the physical link to the spot market in which the building participates. This is represented in the grid cost $C_{\textrm{grid}}$ defined in Eq. 3. The grid power $P_{\textrm{grid}}$ is included in the state vector $S_{a,t}$.
The system is composed of a 5kWp SPV, a 20kWh BESS with nickel manganese cobalt oxides (NMC) or Lithium iron phosphate (LFP) cells, two 12.5kW EV charging points, a 4kWe heat pump, a 2.7kWth solar thermal collector, a 200kWh TESS, a 6kWp electrical load, a 1kWp thermal demand, and 10kW LV grid connection. Power consumption profiles ($P_{\textrm{load}}^{\textrm{e}}$) were constructed for a year using data from 2021 to 2023 from the TU Delft’s Green Village smart meter data [49]. The output of the SPV is taken from [50, 51], the market prices $\lambda$ are taken from the EPEX day-ahead auction, and $\lambda_{\textrm{buy}}=0.95\lambda_{\textrm{sell}}$ [52], and the heat demand $P_{\textrm{load}}^{\textrm{th}}$ was modeled as [30].
The cells used are SANYO NCR18650 cells for NMC as in [15] and A123 cells for LFP. Their datasets where taken from PyBaMM [53] and LiiBRA [46]. To construct the ECM for both cell types sythetic/simulated cells were built and simulated in PyBaMM using standard 1D full order model (FOM). The simulations followed the testing methodology from Plett, Chapter 2 Sections 2.9-.11 , [17]. Once the simulated profiles are ready the ECM parameters can be identified using subspace system identification as in [54]. The parameters can be found in A, Table 5. The capacity fade cost is assumed to be $c_{\text{loss}}=1.2$/Ah, roughly 280-310€/kWh depending on the average voltage.
@@ 5.1 Case Study I: Day-ahead planning
To test and validate our EMA the planner was simulated for two standard months (summer and winter) using 2023 data from the previously mentioned sources.
To quantify the impact of each performance and aging model 3 day-ahead planners were built:
@@
    (CEmpDeg) Including a first-order ECM and empirical aging for the $b$, with $\tilde{S}^{M2}_{a,t}$ in Eq. 10.
@@
    (CPBDeg) Including first-order ECM with PBROM aging for the $b$, with $\tilde{S}^{M3}_{a,t}$ in Eq. 10.
The simulation workflow is presented in Algorithm 2 and depicted in Fig. 2. First, the hyperparameters are initialized. This includes the time window to be optimized $\mathcal{D}_{t}^{\pi}=[t;t+H]$, number of days $n_{d}$, user preferences, the initial state $S_{a,0}$, and weights $w$. In our case for day-ahead planning, the time window is $H=48\text{hr}-\Delta t$ and $n_{d}=29$. The weights are $w_{\text{grid}}=w_{\text{loss}}=1$ and $w_{\text{SoCDep}}=w_{\text{TESS}}=1000$ for the penalty terms. At timestep $t$ the OCP in Eq. 10, is solved obtaining the actions sequence $\mathcal{P}_{a,[t,t+(H+\Delta t)/2]}^{*}$. Together with the exogenous information sequence $\mathcal{W}_{[t+1,t+(H+\Delta t)/2]}$ the actions are passed to the simulator $S_{a,t}^{M}$ to get the feedback state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$. This feedback loop is repeated $n_{d}$ times.
@@
As a representative example Fig. 6 presents the results for the CPBDeg planner for a monthly period. It has the resulting power balances (electrical and thermal) and the use of the hybrid energy storage system (HESS). The electric ESS have daily cycles to minimize operating costs (energy arbitrage). This is particularly important for the EV since its mobility demand already establishes a daily periodicity. Thus, due to the EV’s battery pack size and its natural periodicity, it becomes the main electric storage of the system. This frees up the BESS for energy arbitrage, trying to capture prices variations when possible within the power balance. Thus price volatility incentivizes cycling. However, due to the SEI model, batteries are also pushed downwards to the minimum aging state at $\underline{SoC}_{b,t}$. Thus the dispatch contemplates a trade-off between the 2 parts of the objective $C_{\textrm{grid}}$ and $Q_{\textrm{loss}}$. During the summer, price volatility is high with several hours with $\lambda_{t}\leq 0$, the EMS buys this energy to reduce costs. In winter, prices are less volatile and the load is higher leading to less opportunities for arbitrage and overall higher costs.
Figure 8: Monthly simulation $SoC_{\textrm{EV}}$ for summer (top) and winter (bottom).
On the thermal side, the natural periodicity of the carrier is longer. Thus during the first week the initial $SoC_{\textrm{TESS},0}$ influences costs greatly. After the first week the EMS has already steered the buffer to its desired setpoint. This means a high $SoC_{\textrm{TESS},t}$ during summer (high heat generation, low load) and a low setpoint during winter (low heat generation, high load). The high setpoint during summer entails 2 risks: overcharging the TESS (i.e. activating soft-constraint) and not capturing negative prices/highly volatile prices due to past short-sight (TESS starting a day with a high $SoC$).
The schedules of the HESS under the different planners are summarized in Figs. 7- 9. In the BESS, Fig. 7, the BNoDeg and CEmpDeg cycle the battery pack more often. This is because BNoDeg does not contemplate ageing and the CEmpDeg equations do not relate min $Q_{\text{loss}}$ to $\underline{SoC}$. The CPBDeg planner cycles less frequently, concentrating the operation near $P=0$ and $\underline{SoC}$ to reduce the ageing of the BESS. This is true for both summer (top) and winter (bottom). Moreover, in many days the price variations are not large enough to afford ageing the BESS, thus CPBDeg chooses to maintain the $SoC$ below the benchmarks. The two highlighted days are days were the CPBDeg outperforms the benchmarks and the BESS prioritizes $C_{\text{grid}}$ over $Q_{\text{loss}}$.

### v3

@@ 3.2 Thermal modelling
where $\eta$ denotes a conversion factor or efficiency, $Q_{\textrm{TESS}}$ is the capacity in kWh. To avoid imposing an arbitrary periodicity on the TESS, no terminal conditions are used on its $SoC_{\text{TESS},t}$. Thus, the TESS operates under its natural time constant.
@@ 3.3 Batteries
The remaining devices in the MCES are all battery-based ESS. Batteries have complex nonlinear dynamics, and several modeling techniques are presented in the literature [17]. In this work, models coming from empirical and physics-based approaches are used. The modeling is divided into two different sub-models: performance and ageing. Under the UMF, this is represented in the transition function $\tilde{S}^{M}_{b,t}(\tilde{S}_{b,t},x_{b,t}|\theta_{b,t})$, which contains both the perf. model $p_{b,t}^{M}(.)$ and the ageing model $d_{b,t}^{M}(.)$. The performance model predicts stored energy $SoC_{b,t}$ and terminal voltage $v_{t,b,t}$. The ageing model is used to update the parameters of $p^{M}_{b,t}(.)$, as shown in Fig. 3. Even though the change in parameters $\theta_{b,t}$ becomes significant after considerable ageing has occurred, optimizing it in the short term can lead to considerable savings in the long and medium term [46, 47, 15]. This is because in PB ageing models, the relationship between ageing states and control actions is explicit, and the policy can directly minimize it through the internal states $SoC_{b,t}$.
@@
The ageing model $d_{b,t}^{M}(.)$ is a set of equations that describes the dynamics of the performance parameters $\theta_{b,t}$.
@@
$SoC_{\textrm{BESS},t^{\prime}_{1}}=SoC_{\textrm{BESS},t^{\prime}_{1}+24\text{hs}}$ (20)
where $t_{0}\leq t_{1}\leq H\leq T$. This way, the OCP is better computationally conditioned, but the planner still has the freedom to decide $SoC_{\textrm{BESS},24\textrm{hs}}$. The proposed terminal condition has two key properties: it is more flexible than fixing $SoC_{\text{BESS},24\text{hs}}=50\%$ , and it bounds the value function $V_{H}$ of the OCP. Ideally, no terminal condition would be used to freely use all 3 storage systems. Unfortunately, to solve such an unbounded OCP, an optimization horizon $H$ much larger than 48 hours would be required [48, 49].
@@ 3.3.1 Performance models $p_{b,t}^{M}$
For the performance submodel, two alternatives have been implemented: a simple bucket model (BM) and a first-order equivalent circuit model (ECM). A basic BM of the operation of a battery assumes that its output voltage $v_{t}$ is linear with the state of charge $SoC$, assuming no voltage drop. Hence, the only equations of this model are:
$SoC_{b,t+1}=SoC_{b,t}-\frac{\Delta t}{Q_{b,t}.3600}.\eta_{c}.i_{b,t}\,$, (21a)
$i_{b,t}=\frac{P_{b,t}}{v_{t,b,t}.N_{s,b}.N_{p,b}}\,$, (21b)
$OCV_{b,t}=a_{OCV,b}+b_{OCV,b}.SoC_{b,t}\,$, (21c)
$v_{t,b,t}=OCV_{b,t}\,$, (21d)
$S_{b,t}=[SoC_{b},v_{t,b},i_{b}]^{T}_{t}$ (21e)
where $i_{b,t}$ is the current passing through the cell, $OCV_{b,t}$ is the open circuit voltage, $\eta_{c}$ is the Coulombic efficiency [18] , and $Q_{b,t}$ is the cell capacity in Ah. Each battery pack is assumed to be organized as a series connected module (SCM) where $N_{s/p,\ b}$ are the series cells per branch and parallel branches, respectively. In this model, the most relevant parameter in $\theta_{b}$ is the $Q_{b}$.
Figure 4: First-order Equivalent Circuit Model.
A first-order ECM has improved accuracy due to the incorporation of diffusion and series resistance, as in Fig. 4. The performance sub-model $p_{b,t}^{M}(.)$ is then modified by adding the equation:
@@
where $i_{R_{1},b,t}$ is the current flowing through $R_{1}$ in Fig. 4. Eqs. 21a, 21b and 21c are maintained. The ECM incorporates the series voltage drop that limits power output and the first-order diffusion dynamics. Here, the relevant parameters are $\theta=[Q,R_{0}]^{T}$ that usually define the cell’s state of health $SoH$.
@@
For the ageing models, the first alternative is an empirical sub-model presented by [32]. The empirical sub-model reduces all the degradation mechanisms into calendar and cyclic ageing.
@@
where the fitting parameters $c_{1,...,5}$ are taken from [50, 10, 32] and $t_{0,b}$ is the elapsed lifetime of the battery $b$.
For the physics-based alternative, the reduced order model (PBROM) from [15] is used. It accounts for two degradation mechanisms: the solid electrolyte interface (SEI) and active material loss (AM). The author also presents a PBROM for Li-plating, but given the low C-rate and standard temperature range of this application, it will not be included.
@@
$\eta_{k,b,t}=\frac{2.R.T}{F}.\text{sinh}^{-1}\left(\frac{i_{b,t}}{n_{\textrm{SEI}}.a_{s}.A.L_{n}.i_{0}}\right)$ (27a)
$z_{b,t}=SoC_{b,t}.(z_{100\%,b}-z_{0\%})+z_{0\%,b}$ (27b)
$\beta_{b}=e^{\frac{n_{\textrm{SEI}}.F}{R.T}.\left(\eta_{k,b}+OCV_{n,b,t}-OCV_{s,b}\right)}$ (27c)
where $\eta_{k}$ is the SEI side reaction kinetic overpotential, $z$ is the Li stochiometry of the cell, $OCV_{n}$ is the open-circuit voltage of the anode made with an empirical fit, $OCV_{s}=0.4V$ is the side reaction open-circuit voltage, and $T$ is the cell temperature. It is assumed that the temperature $T$ is constant over time and is controlled by the local primary control system. For the BESS this assumption could be challenged given that the BESS is owned/operated by the same house owners. In the case of the EV, the ownership of the car battery pack sometimes corresponds to the OEM/car manufacturer, limiting access and control over it. The rest of the parameters can be found in the A.
@@
The total ageing is the contribution of both mechanisms SEI layer growth and AM loss. The capacity fade current is:
@@
which is later used again in 25.
@@
To model the power fade (i.e., the increase of $R_{0}$), the SEI layer thickness $\delta_{\textrm{SEI},b,t}$ growth is described by:
@@
Hence, the dynamics of the series resistance $R_{0}$ are:
@@
The solvent S leaves the electrolyte to form the SEI layer; thus, the volume fraction of S evolves with:
@@ 3.4 Electric Vehicle
where $P_{\text{tot},\text{EV},t}$ is the total power of the EV, $P_{\text{EV},t}$ is the charger power, and $P_{\text{drive},\text{EV}}$ is the power consumed driving, assuming no public charging. The total power $P_{\text{tot},\text{EV},t}$ is then used in Eq. (21b) and later for calculating the ageing of the EV batteries. The average driving power is also sampled from a Gaussian distribution $P_{\text{drive},\text{EV}}\sim\mathcal{N}(\mu_{\text{drive}},\,\sigma_{\text{drive}}^{2})$. This is because the EV battery pack degradation during driving needs to be accounted for in the operation strategy (charging and driving).
@@
$\varepsilon_{\textrm{SoC},t_{\text{dep}}}=SoC_{\textrm{EV}}(t_{\textrm{dep}})-SoC_{\textrm{dep}}^{*}$ (36)
Summing up, the three different storage systems have three different ways of handling their terminal set and value function bounds. The BESS has a periodic condition in Eq. 20 , which is a flexible alternative to the classic terminal daily periodicity. The EV follows a soft-tracking problem at $t_{\text{dep}}$. The TESS has no terminal condition. This flexible design is the key contribution to integrating all the storage systems and the non-convex PB models to OCP in Eq. 8.
@@
1:Define setpoint sequences $\mathcal{P}_{a,[t,t+24\text{hs}]}$
2:Define exogenous information sequences $\mathcal{W}_{[t+1,t+1+24\text{hs}]}$
3:Recalculate $P_{\textrm{TESS},t}$ and $P_{\textrm{grid},t}$ with Eqs. 13 , 14, $\mathcal{P}_{a,[t,t+24\text{hs}]}$, and $\mathcal{W}_{a,[t+1,t+1+24\text{hs}]}$
@@
The simulator evaluates the policy $\pi$ and closes the loop with the state measurements. It is designed to:
@@
The whole process is defined in Algorithm 1. First, the power setpoints must be adjusted for the grid and TESS because the forecast used in $X_{t}^{\pi}$ will never be the same as the actual exogenous inputs. Take a look at the balances, Eq. 13 and 14, which contain the loads and solar generation. It is clear that $\tilde{P}\neq P$ and a device must compensate for that difference. Thus, the simulator $S_{a,t}^{M}$ recalculates:
@@
Second, once these powers have been adjusted, the simulator uses these powers to obtain the true/actual/fidelity state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$. For the TESS it recalculates Eq. 17. For the $b$, it uses LiiBRA.jl [52] to swiftly simulate PBROM s of the performance of the battery [17, 53]. After that, the models from Jin [15] are used to calculate the true degradation outcome of the decisions $P_{a,t}^{*}$. Again, the reader must remember that the capacity fade (decrease in $Q_{b,t}$) is modeled in both the simulator $S^{M}_{a,t}$ and the approximate model of the planner $\tilde{S}_{a,t}^{M}$, whereas the power fade (increase in $R_{0,as,t}$) is only addressed in the simulator $S^{M}_{a,t}$. Finally, if an action $P_{b,t}^{*}$ causes a future state to go out of bounds ($S_{b,t+1}\leq\underline{S}_{b}$ or $S_{b,t+1}\geq\overline{S}_{b}$), the remaining actions are rejected and the $b$ remains in that state (either $\overline{S}_{b}$ or $\underline{S}_{b}$) until the next day. Finally, the carriers are rebalanced if necessary.
@@ 5 Case studies
5:  Simulate $S_{a,t+1}=S_{a,t}^{M}(S_{a,t},P^{*}_{a,t},W_{t+1})\ \forall t\in\mathcal{D}_{t}=[t;t+24\text{hs}]$, using Algorithm 1
@@
7:  Move time window $\mathcal{D}^{\pi}_{t}\leftarrow\mathcal{D}^{\pi}_{t}+24\text{hs}$;
@@
The building has a grid connection with a smart meter with 15min resolution. The grid power $P_{\textrm{grid}}$ is included in the state vector $S_{a,t}$.
The system is composed of a 5kWp SPV, a 20kWh BESS with nickel manganese cobalt oxides (NMC) or Lithium iron phosphate (LFP) cells, two 12.5kW EV charging points, a 4kWe heat pump, a 2.7kWth solar thermal collector, a 200kWh TESS, a 6kWp electrical load, a 1kWp thermal demand, and 10kW LV grid connection. Power consumption profiles ($P_{\textrm{load}}^{\textrm{e}}$) were constructed for a year using data from 2021 to 2023 from TU Delft’s Green Village smart meter data [55]. The output of the SPV is taken from [56, 57], the market prices $\lambda$ are taken from the EPEX day-ahead auction, and $\lambda_{\textrm{buy}}=0.95\lambda_{\textrm{sell}}$ [58], and the heat demand $P_{\textrm{load}}^{\textrm{th}}$ was modeled as [34].
The cells used are SANYO NCR18650 cells for NMC as in [15] and A123 cells for LFP [31]. Their datasets were taken from PyBaMM [59] and LiiBRA [52]. The ECM s for both cell types were fitted from synthetic cells simulated in PyBaMM, following the methodology in [17]. Once the simulated profiles are ready, the ECM parameters can be identified using subspace system identification as in [60]. The parameters can be found in A, Tables 4 and 5. The capacity fade cost is assumed to be $c_{\text{loss}}=1.2$ €/Ah, roughly 280-310€/kWh, depending on the average voltage.
@@ 5.1 Case Study I: Day-ahead planning
To test and validate our EMA the planner was simulated for two standard months (summer and winter) using 2023 data from the previously mentioned sources. Both are presented in Fig. 5. To quantify the impact of each performance and ageing model 3 day-ahead planners, 2 benchmarks and our proposed approach, were implemented:
@@
    (CEmpDeg) Including a first-order ECM and empirical ageing for the $b$, with $\tilde{S}^{M2}_{a,t}$ in Eq. 8.
@@
    (CPBDeg) Including first-order ECM with PBROM ageing for the $b$, with $\tilde{S}^{M3}_{a,t}$ in Eq. 8.
The simulation workflow is presented in Algorithm 2 and depicted in Fig. 2. First, the hyperparameters are initialized. This includes the time window to be optimized $\mathcal{D}_{t}^{\pi}=[t;t+H]$, the number of days $n_{d}$, user preferences, the initial state $S_{a,0}$, and weights $w$. In our case for day-ahead planning, the time window is $H=48\text{hr}-\Delta t$ and $n_{d}=29$. The weights are $w_{\text{grid}}=w_{\text{loss}}=1$ and $w_{\text{SoCDep}}=w_{\text{TESS}}=1000$, since the first two represent real economic costs and the second represent penalties. At timestep $t$ , the OCP in Eq. 8, is solved, obtaining the action sequence $\mathcal{P}_{a,[t,t+24\text{hs}]}^{*}$. Together with the exogenous information sequence $\mathcal{W}_{[t+1,t+24\text{hs}]}$ the actions are passed to the simulator $S_{a,t}^{M}$ to get the feedback state sequence $\mathcal{S}_{a,[t,t+24\text{hs}]}$. This feedback loop is repeated $n_{d}$ times.
As a representative example, Fig. 6 presents the results for the proposed CPBDeg planner for a monthly period. It has the resulting power balances (electrical and thermal) and the use of the hybrid energy storage system (HESS). The electric ESS have daily cycles to minimize operating costs (energy arbitrage). This is particularly important for the EV since its mobility demand already establishes a daily periodicity. Thus, due to the EV’s battery pack size and its natural periodicity, it becomes the main electric storage of the system. This frees up the BESS for energy arbitrage, trying to capture price variations when possible within the power balance. Thus, price volatility incentivizes cycling. However, due to the SEI model, batteries are also pushed downwards to the minimum ageing state at $\underline{SoC}_{b,t}$. Hence, the dispatch contemplates a trade-off between the 2 parts of the objective $C_{\textrm{grid}}$ and $Q_{\textrm{loss}}$. On the thermal carrier, the $P_{\text{load},t}^{\text{th}}$ represents the heat losses of the building which have to be compensated by either ST, HP or TESS. In the summer, the losses during solar hours are almost zero and thus the ST charges the TESS. If the prices are negative enough then HP charges the TESS during $\lambda_{t}\leq 0$. In the winter, the TESS is charged with the HP early in the day when prices are low or during solar hours, in that order of priority, to later deliver heat to the demand at high price hours. In summary, during the summer price volatility is high with several hours with $\lambda_{t}\leq 0$, the EMS buys this energy to reduce costs. In winter, prices are less volatile and the load is higher, leading to fewer opportunities for arbitrage and overall higher costs.
The schedules of the HESS under the different planners are summarized in Figs. 7- 9. In the BESS, Fig. 7, the BNoDeg and CEmpDeg cycle the battery pack more often. This is because BNoDeg does not contemplate ageing, and the CEmpDeg equations do not relate min $Q_{\text{loss}}$ to $\underline{SoC}$. The CPBDeg planner cycles less frequently, concentrating the operation near $P=0$ and $\underline{SoC}$ to reduce the ageing of the BESS. This is true for both summer (top) and winter (bottom). Moreover, on many days the price variations are not large enough to afford ageing the BESS, thus CPBDeg chooses to maintain the $SoC$ below the benchmarks. The two highlighted days are days where the CPBDeg outperforms the benchmarks and the BESS prioritizes $C_{\text{grid}}$ over $Q_{\text{loss}}$. Finally the linear BNoDeg is the planner with the highest percentage of rejects, roughly 45% of the total time in the summer. This is because when the BNoDeg plans for high-power discharges its predicted $\tilde{SoC}_{b,t}$ deviates from the true $SoC_{b,t}$ of the simulator $S^{M}_{a,t}$, wrongly depleting the ESS early (lower bound $\underline{SoC}_{b,t}$).
@@
For the EV timeseries comparison, Fig.8, the user’s mobility requirement leads to similar timeseries for all the planners. In general, CPBDeg delays the charging to extend the V2G as much as possible, to avoid $\overline{SoC}_{\textrm{EV}}$. V2G mode is less frequent in the BNoDeg/CEmpDeg than in the CPBDeg because of this, and is valid in both standard seasons. This counterintuitive result is because the ageing PBROM show how $Q_{\text{loss}}$ is minimized by $\underline{SoC}$. Since $p_{\text{SoCDep}}$ requires high $SoC^{*}_{\text{dep}}$, this results in more V2G.
On the thermal side, presented in Fig. 9, the natural periodicity of the carrier is longer due to its thermal efficiencies and size. Hence, during the first week the initial $SoC_{\textrm{TESS},0}$ influences costs greatly. After the first week, the EMS has already steered the buffer to its desired setpoint. This means a high $SoC_{\textrm{TESS},t}$ during summer (high heat generation, low load) and a low setpoint during winter (low heat generation, high load). The high setpoint during summer entails 2 risks: overcharging the TESS (i.e., activating soft-constraint and curtailing ST) and not capturing negative prices prices due to past short-sightedness (TESS starting a day with a high $SoC$). Even though in the summer CPBDeg has the smallest ST curtailment the ageing models have a low overal impact. In other words, integrating ageing models does not change the overall TESS decisions.
Figure 8: Monthly simulation $SoC_{\textrm{EV}}$ for summer (top) and winter (bottom). Grey bands indicate driving periods.
In summary, each storage system follows a distinct use represented in its terminal set. The EV follows a soft-tracking penalty $p_{\text{SoCDep}}$, the BESS follows a flexible daily periodicity, and the TESS handles medium-term thermal demand. Without the design of the mixed terminal set, the non-convex NLP would not converge to feasible locally optimal points, and thus is a key contribution of this paper.